Logiweb dictionary of factorial

Up Help

0 0 factorial

1 1 ∗!

The pyk compiler, version 0.grue.20050502+ by Klaus Grue GRD-2005-05-26.UTC:10:24:54.494911 = MJD-53516.TAI:10:25:26.494911 = LGT-4623819926494911e-6